OpenStudy (anonymous):

Please help me out. :) Two charges, one twice as large as the other, are located 15 cm apart and experience a repulsion force of 95 N. What is the magnitude of the larger charge?

ganeshie8 (ganeshie8):

use coulomb's law \[F = k\dfrac{q_1q_2}{r^2}\]

OpenStudy (anonymous):

ahh.. okay.. I'll try to solve. :)

ganeshie8 (ganeshie8):

just plugin the given values after converting to appropriate units k = 9*10^9 F = 95 r = 15*10^-2

OpenStudy (anonymous):

Q1Q2 = 2.375 x 10^-10 C^2

ganeshie8 (ganeshie8):

looks good!

OpenStudy (anonymous):

Q1 = Q Q2 = 2Q

ganeshie8 (ganeshie8):

yes solve Q, just algebra..

OpenStudy (anonymous):

Ahh.. But I don't know how to separate Q1 and Q2 in the equation.. How will I do that sir? :)

ganeshie8 (ganeshie8):

Notice Q1 = Q2/2

ganeshie8 (ganeshie8):

Q1Q2 = 2.375 x 10^-10 C^2 becomes (Q2/2)Q2 = 2.375 x 10^-10 C^2

OpenStudy (anonymous):

Notice Q1 = Q2/2 .. How come?

OpenStudy (anonymous):

Ahh.. Got it... :) Thanks. :)

ganeshie8 (ganeshie8):

We're given one charge is twice as the other that means the smaller charge is half of the bigger charge, yes ?

ganeshie8 (ganeshie8):

Q1Q2 = 2.375 x 10^-10 C^2 becomes (Q2/2)Q2 = 2.375 x 10^-10 C^2 Q2^2/2 = 2.375 x 10^-10 C^2 Q2^2 = 2*2.375 x 10^-10 C^2 Q2 = sqrt(2*2.375 x 10^-10 )

OpenStudy (anonymous):

Thank you! :) I understand it well. :)